OpenStudy (anonymous):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (amistre64):

they can all be solved by factoring since all quadratics have 2 roots

OpenStudy (amistre64):

either 2 real roots 2 repeated roots or 2 complex roots

OpenStudy (anonymous):

I need a numerical answer*

OpenStudy (amistre64):

do you remember what we just went over in the last question? what part defines how many roots it gets?

OpenStudy (anonymous):

Yes.

OpenStudy (amistre64):

the roots define the factors, so the thing that defines the roots also defines the factoring

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

can you show me which one that is in the options above?

OpenStudy (anonymous):

For the question I have to answer it's fill in the blank so I don't have any choices to chose from.

OpenStudy (amistre64):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b b2 − 4ac 2a ^^^^^^^^^^^^^^^^^^^ these look like choices to me

OpenStudy (anonymous):

Oh whoops. Yes those are. Sorry :/

OpenStudy (amistre64):

since the radicand defines the roots, and the roots define if it can be factored; which option is most like the radicand part of the quadratic formula?

OpenStudy (anonymous):

b^2-4ac?

OpenStudy (amistre64):

exactly

OpenStudy (amistre64):

now i spose they gave you a specific example to pull a b and c from

OpenStudy (anonymous):

yes 2x^2 + 7x + 3 = 0

OpenStudy (amistre64):

can you define a b and c from that for me?

OpenStudy (anonymous):

a= 2 b= 7 and c=3

OpenStudy (amistre64):

correct, so lets fill those in to the discriminant b^2 - 4ac 7^2 -4(2)(3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Once they are filled in what do we do? Is that the final answer?

OpenStudy (anonymous):

if the discriminant is a perfect square, like 25 in this case, you can easily solve by factorizing.

OpenStudy (anonymous):

Sorry this really is not my best math topic. I know how to factor but im not sure what I should be factoring in this problem.

OpenStudy (anonymous):

But i dont think problem wants you to factor anything. it wants you to calculate the part that tells you if it can be solved by factor. and that answer =b^2-4ac= 25. Two things are needed to understand this prob - quadratic forumula and solving quadratic by fatcors . only then it wil make sense

OpenStudy (anonymous):

Ok thank you (: